Find the distance between (7, – 2) and (– 5, 3)
Answer by checkley75(3666)

THE TWO SIDES OF THIS RIGHT TRIANGLE ARE:
7+5=12 & 3=2=5. THEREFORE USING THE PATHOGREAN THEORUM WE GET
12^2+5^2=D^2
144+25=D^2
169=D^2
16=D WHICH IS THE DISTANCE BETWEEN THESE POINTS.